OpenStudy (aravindg):

Note:This is not a question This is my first tutorial and i begin with some maths tricks that i hope you all will like ..do provide your feedback! thanks! P.S. more tutorials coming :)

OpenStudy (aravindg):

Square of numbers near to 100 ------------------------- Let me explain this trick by taking examples 96^2 :- First calculate 100-96, it is 4 so 96^2 = (96-4)----4^2 = 9216 similarly 106^2 :- First calculate 106-100, it is 6 so 106^2 = (106+6)----6^2 = 11236 Now , An other case arises 110^2 = (110+10)----100 = (120+1)----00 = 12100 similarly 89^2 = (89-11)----121 = (78+1)----21 = 7921

OpenStudy (anonymous):

not only for hundred , it is for every no. near 10,100,1000,10000,etc.

OpenStudy (aravindg):

@shubham.bagrecha indeed we can extend to 100,1000,....etc using the similiar method

OpenStudy (anonymous):

Can u explain why it works..

hartnn (hartnn):

even i didn't get this : 96^2 = (96-4)----4^2 = 9216

OpenStudy (anonymous):

96-4=92 4^2=16 So, it becomes 9216

OpenStudy (anonymous):

Like adding strings

OpenStudy (anonymous):